Using the Ideal Gas Law, fill in the missing variable in the table below: P V...

  1. Using the Ideal Gas Law, fill in the missing variable in the table below:

P

V

n

T

a

2.00 L

2.300 mole

127 C

6.00 atm

b

4.00 mole

-73 C

3.75 atm

3.75 L

c

100 C

1.22 atm

2.00 L

0.500 mole

d
Answer #1

a)

Given:

V = 2.0 L

n = 2.300 mol

T = 127.0 oC

= (127.0+273) K

= 400 K

use:

P * V = n*R*T

P * 2 L = 2.3 mol* 0.08206 atm.L/mol.K * 400 K

P = 37.7476 atm

Answer: 37.7 atm

b)

Given:

P = 6.0 atm

n = 4.00 mol

T = -73.0 oC

= (-73.0+273) K

= 200 K

use:

P * V = n*R*T

6 atm * V = 4 mol* 0.08206 atm.L/mol.K * 200 K

V = 10.9413 L

Answer: 10.9 L

c)

Given:

P = 3.75 atm

V = 3.75 L

T = 100.0 oC

= (100.0+273) K

= 373 K

find number of moles using:

P * V = n*R*T

3.75 atm * 3.75 L = n * 0.08206 atm.L/mol.K * 373 K

n = 0.4594 mol

Answer: 0.459 mol

d)

Given:

P = 1.22 atm

V = 2.0 L

n = 0.500 mol

use:

P * V = n*R*T

1.22 atm * 2 L = 0.5 mol* 0.08206 atm.L/mol.K * T

T = 59.5 K

Answer: 59.5 K